# On the existence of unbiased resilient estimators in discrete quantum systems

### 23 Feb 2024 v1

@@ I Introduction
Quantum metrology studies the limits in precision while estimating physical parameters encoded in quantum systems. Quantum properties such as entanglement and squeezing are used to surpass classical estimation strategies [1, 2, 3]. Applications like quantum phase estimation [4, 5, 6, 7], quantum-enhanced position and velocity estimation [8, 9, 10], quantum illumination [11, 12, 13], quantum thermometry [14, 15, 16, 17], and channel discrimination [18] among others have boosted the field. Parameter estimation gives the theoretical framework to establish the optimal performance of particular estimation protocols and its main figure of merit is the Cramér-Rao bound (CRB) [19, 20].
@@ II Estimation theory
To extract information from a physical system measurements have to be performed. The measurement outcome is modeled as a random variable following a probability distribution that depends on the state of the system. The information is encoded in the random variable and estimation theory is the branch of statistics that deals with the optimal protocols and the ultimate limit for the information extraction. The CRB establishes a limit on the precision while estimating a magnitude or a parameter. Consider that the outcome of the experiment, denoted by the random variable $X$, follows the family of probability distribution $P_{\bm{\theta}}(x)$, that is parameterized by the vector $\bm{\theta}=(\theta_{1},\theta_{2},...)$ which is being estimated. In this article, we only focus on single parameter estimation, i.e. $\bm{\theta}=\theta$. An estimator $\tilde{\Theta}(x)$ is a function of the random measure outcome that estimates the value of the parameter $\theta$. The performance of an estimator is characterized by its bias $b(\theta)$, mean squared error (MSE) and variance $(\Delta\tilde{\Theta})^{2}$,
@@ III.1 Classical
$A=\begin{pmatrix}P_{\theta_{0}}(x_{1})&P_{\theta_{0}}(x_{2})&\dots&1-\sum_{i}^{N-1}P_{\theta_{0}}(x_{i})\\ \partial P_{\theta_{0}}(x_{1})&\partial P_{\theta_{0}}(x_{2})&\dots&-\sum_{i}^{N-1}\partial P_{\theta_{0}}(x_{i})\\ \partial^{2}P_{\theta_{0}}(x_{1})&\partial^{2}P_{\theta_{0}}(x_{2})&\dots&-\sum_{i}^{N-1}\partial^{2}P_{\theta_{0}}(x_{i})\\ \vdots&\vdots&\ddots&\vdots\\ \partial^{n}P_{\theta_{0}}(x_{1})&\partial^{n}P_{\theta_{0}}(x_{2})&\dots&-\sum_{i}^{N-1}\partial^{n}P_{\theta_{0}}(x_{i})\\ \\ \end{pmatrix},\quad\quad\vec{x}=\begin{pmatrix}\tilde{\theta}_{1}\\ \tilde{\theta}_{2}\\ \vdots\\ \tilde{\theta}_{N}\end{pmatrix},\quad\quad\vec{b}=\begin{pmatrix}\theta_{0}\\ 1\\ 0\\ \vdots\\ 0\end{pmatrix}.$ (24)
The system has $n+1$ equations and $N$ variables. It is also important to stress that $\partial^{l}P_{\theta_{0}}(x)=\frac{\partial^{l}P_{\theta}(x)}{\partial\theta^{l}}|_{\theta=\theta_{0}}$. We consider the case where the first $N-1$ derivatives are linearly independent. This implies that if $n\leq N-1$, there is always a solution. If $n>N-1$, then the higher derivatives are no longer independent, and they can be written as a combination of the first $N-1$ linearly independent derivatives. In this case, the system has no solution or the solution is trivial, that is to say, the $n^{th}$order BhB is equal to $N-1$ order BhB for all $n>N-1$. It is then concluded that given a $N$ points probability distribution is advisable to compute BhB bound up to order $N-1$.

### 01 Mar 2024 v2

@@ I Introduction
Quantum metrology studies the limits in precision by estimating physical parameters encoded in quantum systems. Quantum properties such as entanglement and squeezing are used to surpass classical estimation strategies [1, 2, 3]. Applications like quantum phase estimation [4, 5, 6, 7], quantum-enhanced position and velocity estimation [8, 9, 10], quantum illumination [11, 12, 13], quantum thermometry [14, 15, 16, 17], and channel discrimination [18] among others have boosted the field. Parameter estimation gives the theoretical framework to establish the optimal performance of particular estimation protocols and its main figure of merit is the Cramér-Rao bound (CRB) [19, 20].
@@ II Estimation theory
To extract information from a physical system measurements have to be performed. The measurement outcome is modeled as a random variable following a probability distribution that depends on the state of the system. The information is encoded in the random variable and estimation theory is the branch of statistics that deals with the optimal protocols and the ultimate limit for the information extraction. The CRB establishes a limit on the precision while estimating a magnitude or a parameter. Consider that the outcome of the experiment, denoted by the random variable $X$, follows the family of probability distribution $P_{\theta}(x)$, that is parameterized by the parameter $\theta$ which is being estimated. An estimator $\tilde{\Theta}(x)$ is a function of the random measurement outcome that estimates the value of the parameter $\theta$. The performance of an estimator is characterized by its bias $b(\theta)$, mean squared error (MSE) and variance $(\Delta\tilde{\Theta})^{2}$,
@@ III.1 Classical
$A=\begin{pmatrix}P_{\theta_{0}}(x_{1})&P_{\theta_{0}}(x_{2})&\dots&1-\sum_{i}^{N-1}P_{\theta_{0}}(x_{i})\\ \partial P_{\theta_{0}}(x_{1})&\partial P_{\theta_{0}}(x_{2})&\dots&-\sum_{i}^{N-1}\partial P_{\theta_{0}}(x_{i})\\ \partial^{2}P_{\theta_{0}}(x_{1})&\partial^{2}P_{\theta_{0}}(x_{2})&\dots&-\sum_{i}^{N-1}\partial^{2}P_{\theta_{0}}(x_{i})\\ \vdots&\vdots&\ddots&\vdots\\ \partial^{n}P_{\theta_{0}}(x_{1})&\partial^{n}P_{\theta_{0}}(x_{2})&\dots&-\sum_{i}^{N-1}\partial^{n}P_{\theta_{0}}(x_{i})\\ \\ \end{pmatrix},\quad\quad\vec{x}=\begin{pmatrix}\tilde{\theta}_{1}\\ \tilde{\theta}_{2}\\ \vdots\\ \tilde{\theta}_{N}\end{pmatrix},\quad\quad\vec{b}=\begin{pmatrix}\theta_{0}\\ 1\\ 0\\ \vdots\\ 0\end{pmatrix},$ (24)
where we defined $\partial^{l}P_{\theta_{0}}(x)\equiv\frac{\partial^{l}P_{\theta}(x)}{\partial\theta^{l}}|_{\theta=\theta_{0}}$. The system has $n+1$ equations and $N$ variables. We consider the case where the first $N-1$ derivatives are linearly independent. This implies that if $n\leq N-1$, there is always a solution. If $n>N-1$, then the higher derivatives are no longer independent, and they can be written as a combination of the first $N-1$ linearly independent derivatives. In this case, the system has no solution or the solution is trivial, that is to say, the $n^{th}$order BhB is equal to $N-1$ order BhB for all $n>N-1$. It is then concluded that given a $N$ points probability distribution is advisable to compute BhB bound up to order $N-1$.
@@ III.2 Quantum
Let us elaborate more on this because at first glance the result may seem to be at odds with the classical case. After projective measurements, one deals with $N$ point probability distributions. Therefore, the results derived in the classical setting must also apply, so it might seem that there will be a contradiction to what we stated here. To solve this apparent contradiction, we explain why $N$ order QBh bound need not be equal to the $N-1$ one. The key point is that to compute the $N$ and the $N+1$ order QBhB, different measurements have to be performed, giving different probability distributions $P_{N}$ and $P_{N+1}$. This extra freedom in the measurements is what gives the quantum systems more non-trivial information, or in other words, the maximal order for which QBhB gives different information is higher than for classical scenarios. Thus, since $P_{N}$ and $P_{N+1}$ are different probability distributions, there is no discrepancy with what we presented in section (III.1).